[Q19-20] You are the CFO of a major pharmaceutical firm. A division manager has presented senior management with an investmen Need help with 19 & 20
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

Since it is mentioned the project is independent of the outcome of general market

=> beta = 0

r=rt + B* rm -rf)

=> r = 3% + 0*(12%-3%) = 3%

Expected value of Project in one year = $1 billions * .1 + $*0.9

Expected value of Project in one year = $100 millions

NPV = Exp Expected value of Project in one year (1 + 0.03) - Initial cost

100 NPV = (1 + 0.03) - 95

NPV = 97.1 - 95

NPV = 2.1 million